OpenStudy (anonymous):

HELP PLEASE!!!! A MEDAL FOR WHOEVER GETS IT RIGHT!!! Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (theraggedydoctor):

divide both sides by -3

OpenStudy (theraggedydoctor):

|x + 7| = 4

OpenStudy (anonymous):

ok

OpenStudy (theraggedydoctor):

x + 7 = 4 x - 7 = -4 x = -3 x = 3

OpenStudy (anonymous):

then do i subtract 7?

OpenStudy (mathmale):

Excuse my saying this, but the goal of Open Study is to help YOU "GET IT RIGHT." Not to wait for others to do the work for you and provide you with the answers. Because your math problem involves the absolute value operator ( | | ), we need to solve this equation twice! -3|x + 7| = -12 has two cases: 1) x+7 is already positive, so the absolute value operator is not needed. -3(x + 7) = -12 Please solve this for x. 2) x+7 is negative, so the absolute value operator makes it positive. -3[-(x + 7)] = -12 Please solve this for x. You'll end up with 2 different answers. Be sure to check both by substituting them into the original equation. Your turn.

OpenStudy (mathmale):

I'd encourage you to share your work as you do it. Then others could give you specific feedback on your work. If you're working on -3(x + 7) = -12, multiply each of the 2 terms inside parentheses by -3. Then think about what you 'll need to do to solve for x.

OpenStudy (mathmale):

TheRaggedyDoctor has already given you a very good suggestion: "divide both sides by -3"

OpenStudy (mathmale):

Divide both sides by -3: -3(x + 7) = -12 This is a better approach than the one I suggested earlier.